OpenStudy (anonymous):

Determine the volume of the region that lies under the sphere x^2+y^2+z^2=9 , above the plane z=0 and inside the cylinder x^2+y^2=5.

OpenStudy (anonymous):

Using polar coordinates btw.

OpenStudy (anonymous):

So i found the limits of the double intergal

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{\sqrt{5}}r \sqrt{9-r^2}drd \theta\]

OpenStudy (anonymous):

But I am just wondering why the equation of the cylinder wasn't incorporated into the integral.

OpenStudy (anonymous):

@dan815 @zzr0ck3r

OpenStudy (anonymous):

NVM!!! :P . Got it >.< .